# Modeling error correction with Lindblad dynamics and approximate channels

### v2

@@ Abstract
We analyze the performance of a quantum error correction code subject to physically motivated noise modeled by a Lindblad master equation. We consider dissipative and coherent single-qubit terms and two-qubit crosstalk, studying how different approximations of the noise capture the success rate of a code. Focusing on the five-qubit code and adapting it to partially correct two-qubit errors in relevant parameter regimes according to the noise model, we find that a composite-channel approximation where every noise term is considered separately captures the behavior in many physical cases up to long timescales, eventually failing due to the effect of noncommuting terms. In contrast, we find that single-qubit approximations do not properly capture the error correction dynamics with two-qubit noise, even for short times. A Pauli approximation going beyond a single-qubit channel is sensitive to the details of the noise, state, and decoder, and succeeds at short timescales relative to the noise strength, beyond which it fails. Furthermore, we point out a mechanism for a Pauli model failure where it underestimates the success rate of a code even with frequent syndrome projection and correction cycles. These results shed light on the performance of error correction in the presence of realistic noise and can advance the ongoing efforts towards useful quantum error correction.
@@ I Introduction
However, Pauli noise channels are just a subset of completely positive and trace preserving (CPTP) quantum operations (or Kraus channels). Moreover, multiqubit noise channels can rarely be derived explicitly in physically-relevant cases [17]. In contrast, the generators of continuous dynamics of individual single qubits (1Q) and interacting two-qubit (2Q) pairs are often simpler to handle. Using the generators, the evolution of open systems is described by a master equation, and when the environment is memoryless (the dynamics being Markovian), the Lindblad master equation is the standard choice since it guarantees CPTP dynamics by construction [18, 19]. Lindbladian systems are extensively studied, with analytic, numerical and approximate solution methods continuously being improved [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. Nevertheless, only a few existing works have compared error correction success rates between physically-motivated noise descriptions and their Pauli approximations. The majority of previous work considered only 1Q noise terms [33, 34, 35, 36, 37]. Other works have focused on the effect of coherent errors (often considering analytically solvable cases), as in [38, 39, 40, 41], and crosstalk and gate errors have been simulated using Kraus channels with small codes [42, 5], or twirled to obtain Pauli channels as in [43]. Some works devoted to linking a basic Lindblad model of a system of qubits to error correction success rates include [44, 45] and some comparison to experiment in [46].
In this paper, we study the accuracy of different approximations in modelling the behavior of a quantum error correction code when considering hardware-relevant Lindblad noise. We consider dynamics generated by 1Q coherent (unitary) and incoherent (dissipative) noise and also 2Q coherent interactions. We focus on the success rate of the five-qubit code [47], a toy-model code capable of correcting both bit- and phase-flip errors [40, 48], which we adapt to decode and correct 2Q errors (at the expense of some of the 1Q errors), according to the noise model. We study the question of quantum memory decoherence, with a logical code state evolving subject to the noise for varying durations. We compare the performance of the code with the state evolved under several approximations: (i) continuous Lindblad dynamics, the closest approach to realistic dynamics among those considered here (and the most computationally demanding one), (ii) an intermediate approach consisting of a composition of partial 2Q and 1Q Kraus channels, (iii) a Pauli approximation of the channels, and (iv) an attempt at an approximation using 1Q channels only.
@@
The main results of this paper are as follows. The composite-channel approximation, which is based on noise generators but does not take into account the non-commutativity between noise terms, is a good approximation for relevant ranges of noise parameters, and up to considerably long time scales. We point out the regime of its failure due to the neglect of noncommuting error terms. On the other hand, the validity of the multi-qubit Pauli approximation strongly depends on the noise parameters, initial state and decoder. For many cases it holds pretty well for short time scales, which are most relevant for error correction, and diverges significantly for longer times, and our study exemplifies a number of caveats of the Pauli approximation, whose relevance is broad and extends beyond the code capacity model. In particular, the Pauli noise model could fail in capturing effects of coherent noise and accumulation of phases in logical errors, even when frequently repeated syndrome cycles are applied. Therefore, before using this approximation it would seem necessary to check its validity for the specific details of the studied setup.
@@ II The Lindbladian noise model
where $\sigma^{a}_{i}$ is the Pauli matrix acting on qubit $i$ with $a\in\{x,y,z\}$, $I_{i}$ is the single-qubit identity matrix, the sum $\sum_{i}$ runs over all the qubits, $\sum_{\left\langle ij\right\rangle}$ runs over all coupled pairs, and $h_{i}$, $\zeta_{i,j}$ are coefficients in units of angular frequency. $H_{i}$ corresponds to a discrepancy between the actual qubit frequency and our estimated value of it (determined by the precision of the calibration experiment and ensuing drifts), and $V^{zz}_{ij}$ is a residual crosstalk between coupled qubits, resulting from uncontrolled interactions of ZZ type (also referred to as Ising coupling). In physical systems the crosstalk would often couple nearest-neighbor qubits that would typically be data and parity (auxiliary) qubits, and parity qubits are lacking in our model to begin with. Nevertheless, effective longer-range interaction is also possible, and we consider this coupling as a representative term of coherent dynamics between qubits that are coupled either directly or indirectly. Some 2Q coherent dynamics is also expected when taking into account the over/under rotations of realistic 2Q gates.
@@ III.1 Estimating error correction success rates
The next question is how to extract a useful metric from $\rho(t)$. As discussed further below, an explicit analysis that accounts for all possible syndromes and their corresponding recoveries could become challenging with large codes due to the exponential scaling of their number. Instead, we present in this subsection a simple and efficient approach to compute the code performance, suitable within the code capacity model and scalable to larger codes.
Every error correction code has a set of correctable errors that the code can rectify, $C\equiv\left\{C_{j}\right\}$. These errors can be chosen in such a way that their application to a logical code state results in orthogonal states, as discussed in [47]. Starting with an initial ideal logical state $|\psi_{0}\rangle$, we define $|\varphi_{j}\rangle=C_{j}|\psi_{0}\rangle$. If following the noise operation (before the syndrome extraction and recovery process) the system is in one of the states $|\varphi_{j}\rangle$ then the error correction procedure will successfully recover the initial state. If the system is in any other state, and since we are in the code capacity model, the result will be a different logical state, which will be orthogonal to the initial logical state.
Therefore, for a given state $\rho$, the probability of success is given by
@@
It is worth highlighting the distinction between this method and the conventional approach for assessing the success rate of a code (based on a stochastic Pauli model) presented in Sec. I, in which every time a logical operation is applied is categorized as a failure. The approach that we take here depends on the initial state; if the error correction procedure results in a logical operation that does not alter the initial state, for example, a logical $Z_{L}$ applied to the state $|0\rangle_{L}$, it is considered a success. In this sense, the conventional approach is a worst-case approach deeming every logical operator as a logical error, and this may lead to slight discrepancies between the two methods when applied on the same Pauli noise.
@@ III.2 Estimating a pseudo-threshold
A quantum error correction code is advantageous compared to a physical qubit only for noise parameters such that the probability of an error on the physical qubit is larger than the probability of a logical error on the logical qubit with error correction. Assuming a monotonous dependence on one common noise parameter (where the logical qubit is advantageous for small enough values of the noise parameter), the value of this parameter for which the physical and logical error probabilities are equal determines the pseudo-threshold [34, 56].
@@
We note that the standard way to compute the pseudo-threshold must account for the noise affecting logical states during the preparation, stabilizer measurements and corrections. However, here we are within the simplified framework where all of these are noiseless, giving the logical qubit an unfair advantage, so the pseudo-threshold we obtain is merely a code-capacity pseudo-threshold. Moreover, the 2Q crosstalk does not affect a single (isolated) physical qubit when standing idle, so defining the pseudo-threshold on a single qubit possibly gives the physical qubit an unfair advantage.
@@ III.3 The logical density matrix
To summarize, $\alpha$ is the infidelity of the corrected state with the ideal logical state, or in other words the probability of failure of the correction, and $\beta$ is a measure of coherent logical error that accumulates within the code space. Note that $\eta$ which was studied extensively above, is an approximation of $\alpha$ without explicitly performing the corrections in the circuit. Within the code-capacity model, when the projection onto the syndrome subspaces and the correction are both ideal, $\rho^{P}$ is guaranteed to be within the code space.
@@ IV.3 The Pauli approximation
The main advantage of a Pauli channel is that it allows one to compute the syndromes without simulating the state evolution but rather with the parity matrix, as explained above. The main physically-driven motivations behind this approximation stems from the influence of the syndrome measurement on the noise channel, which reduces many non-Pauli terms, explored in more detail in App. F.
The most precise way to perform the Pauli approximation is by approximating the whole-system quantum channel. However, this requires constructing the channel which may be numerically as challenging as simulating its action on particular state (or even more). Here we choose a less exact but easier to compute approximation, which consists of approximating each individual channel separately. The channels $\mathcal{K}_{\text{1Q}}$ and $\mathcal{U}^{zz}_{\text{2Q}}$ are each approximated by 1Q and 2Q Pauli channels, respectively, which are then composed. A more detailed explanation on the different possible ways to perform this Pauli approximation is given in App. G. We note that the alternative method of approximating the 1Q and 2Q channels together into one 2Q Pauli channel is not a better choice due to the complication arising from the need to “split” the 1Q channels into the different 2Q channels in which each qubit is included, which reduces the accuracy of the approximation.
@@ V.2 The long time limit
$t$ Dynamical Composite, order 1 Composite, order 2 Pauli
@@
Table 2: Table of values of $\eta$ for different approximations and different noise durations. The noise parameters are given in Eqs. (40)-(41). The values presented in this table are $\left\langle\eta\right\rangle\left(\text{std}\left(\eta\right)\right),\sqrt{\left\langle\left(\eta-\eta_{\rm{D}}\right)^{2}\right\rangle}$ where $\eta_{\rm{D}}$ is $\eta$ under the dynamical simulation of the noise, and the average and standard deviation are taken over the six Pauli eigenstates as initial states [Eq. (9)]. In this case the composite-channel approximation does very well, while the Pauli approximation is good for short time scales, but differs significantly for longer time-scales, for which, in this example, it overestimates the failure rate.
For the Kraus channels in the long-time limit, we have $\rho(t\to\infty)\to|0\rangle\langle 0|^{\otimes N}$. However, for the Pauli channels in the long-time limit, for any finite value of $T_{1}<\infty$ we have $\rho(t\to\infty)\to 1\kern-2.5pt\text{l}/2^{N}$, since the 1Q Pauli approximation reduces to the fully depolarizing channel in this limit. This difference in the steady state necessarily implies that at least for $t\gtrsim T_{1}$ (if not earlier), we can a-priori expect any Pauli approximation to break  [^3] The steady state being the fully mixed state results in $\eta\to 1/2$ in the Pauli approximations for any initial state. In contrast, noticing that the state $|0\rangle_{L}$ ($|1\rangle_{L}$) is a superposition over the 5-bit strings with even (odd) number of 1’s, we can calculate the fidelity and correctability of the state in the composite-channel approximation at $t\to\infty$, which gives us $\eta\to 3/8$ for $|0\rangle_{L}$, $\eta\to 5/8$ for $|1\rangle_{L}$ and $\eta\to 1/2$ for $|+\rangle_{L}$..
@@ V.3 Dominant ZZ crosstalk noise
This data strengthens the observations made above, that for the studied parameters the composite-channel approximation is a very good approximation up to intermediate times, while the Pauli approximation works well only for relatively short time-scales. Regarding the composite-channel approximation losing accuracy, this results from the noncommutativity of the 2Q crosstalk channel with amplitude damping. This is a general observation about the nature of the noise model, and the effect of noncommuting terms in the approximation is further explored in App. I.
In the current example, the crosstalk noise is the largest term, and is one that cannot be corrected with the employed standard decoder of the 5Q code. The Pauli approximation overestimates the failure rate when leaving the short time regime, plausibly since it does not describe the coherent dynamics accurately enough. In order to gain more insight into it, we consider in the next subsection the question of a decoder tailored to the noise.
@@ V.4 Modified decoders
In Fig. 3 we present the difference between $\eta$ with the FC ZZ decoder and with the standard decoder, showing that the FC ZZ decoder is beneficial in a wide range of experimentally relevant parameters. Specifically, the parameter values of Eqs. (40)-(41) fall in the regime where the modified decoder is advantageous. Therefore, applying the FC ZZ decoder to the dynamics with the same parameters as in Fig. 2(a), the FC ZZ decoder performs much better than the regular decoder, obtaining much lower values of $\eta$, as can be seen in Fig. 4. We also notice that the Pauli approximation begins to diverge from the dynamical simulation as early as $t\approx 0.7\mu$s, and now it underestimates the failure rate. In other words, even at lower failure rates and when the code is adapted to the dominant noise terms, the Pauli approximation could lose its accuracy already in the short time regime, and not provide even an upper bound on the logical error.

### v3

@@ Abstract
We analyze the performance of a quantum error correction code subject to physically motivated noise modeled by a Lindblad master equation. We consider dissipative and coherent single-qubit terms and two-qubit crosstalk, studying how different approximations of the noise capture the success rate of a code. Focusing on the five-qubit code and adapting it to partially correct two-qubit errors in relevant parameter regimes according to the noise model, we find that a composite-channel approximation where every noise term is considered separately captures the behavior in many physical cases up to long timescales, eventually failing due to the effect of noncommuting terms. In contrast, we find that single-qubit approximations do not properly capture the error correction dynamics with two-qubit noise, even for short times. A Pauli approximation going beyond a single-qubit channel is sensitive to the details of the noise, state, and decoder, and succeeds at short timescales relative to the noise strength, beyond which it fails. Furthermore, we point out a mechanism for a Pauli model failure where it underestimates the failure rate of a code even with frequent syndrome projection and correction cycles. These results shed light on the performance of error correction in the presence of realistic noise and can advance the ongoing efforts towards useful quantum error correction.
@@ I Introduction
However, Pauli noise channels are just a subset of completely positive and trace preserving (CPTP) quantum operations (or Kraus channels). Moreover, multiqubit noise channels can rarely be derived explicitly in physically-relevant cases due the large number of parameters required for the representation. In contrast, the underlying generators of continuous dynamics of individual single qubits (1Q) and interacting two-qubit (2Q) pairs are often simpler to handle and characterize using well-established methods [17, 18]. Using the generators, the evolution of open systems is described by a master equation, and when the environment is memoryless (the dynamics being Markovian), the Lindblad master equation is the standard choice since it guarantees CPTP dynamics by construction [19, 20, 21]. Lindbladian systems are extensively studied, with analytic, numerical and approximate solution methods continuously being improved [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34]. Nevertheless, only a few existing works have compared error correction success rates between physically-motivated noise descriptions and their Pauli approximations. The majority of previous work considered only 1Q noise terms [35, 36, 37, 38, 39]. Other works have focused on the effect of coherent errors (often considering analytically solvable cases), as in [40, 41, 42, 43, 44, 45], and crosstalk and gate errors have been simulated using Kraus channels with small codes [46, 5], or twirled to obtain Pauli channels as in [47]. Some works devoted to linking a basic Lindblad model of a system of qubits to error correction success rates include [48, 49], a comparison to experiment in [50], and numerics going beyond Lindblad dynamics (departing from the Born-Markov approximation for 1Q noise) in [51].
In this paper, we study the accuracy of different approximations in modelling the behavior of a quantum error correction code when considering hardware-relevant Lindblad noise. We consider dynamics generated by 1Q coherent (unitary) and incoherent (dissipative) noise and also 2Q coherent interactions. We focus on the success rate of the five-qubit code [52], a toy-model code capable of correcting both bit- and phase-flip errors [43, 53], which we adapt to decode and correct 2Q errors (at the expense of some of the 1Q errors), according to the noise model. We study the question of quantum memory decoherence, with a logical code state evolving subject to the noise for varying durations. We compare the performance of the code with the state evolving under several approximations of the noise: (i) continuous Lindblad dynamics, the closest approach to realistic dynamics among those considered here (and the most computationally demanding one), (ii) an intermediate approach consisting of a composition of partial 2Q and 1Q Kraus channels, (iii) a Pauli approximation of the channels, and (iv) an attempt at an approximation using 1Q channels only.
@@
The main results of this paper are as follows. The composite-channel approximation, which is based on noise generators but does not take into account the non-commutativity between noise terms, is a good approximation for relevant ranges of noise parameters, and up to considerably long time scales. We point out the regime of its failure due to the neglect of noncommuting error terms. On the other hand, the validity of the multi-qubit Pauli approximation strongly depends on the noise parameters, initial state and decoder. For many cases it holds pretty well for short time scales, which are most relevant for error correction, and diverges significantly for longer times, and our study exemplifies a number of caveats of the Pauli approximation, whose relevance is broad and extends beyond the code capacity model. In particular, the Pauli noise model could fail in capturing effects of coherent noise and accumulation of phases in logical errors. In Sec. V.8 we further show this failure when frequently repeated syndrome cycles are applied.
@@ II The Lindbladian noise model
where $\sigma^{a}_{i}$ is the Pauli matrix acting on qubit $i$ with $a\in\{x,y,z\}$, $I_{i}$ is the single-qubit identity matrix, the sum $\sum_{i}$ runs over all the qubits, $\sum_{\left\langle ij\right\rangle}$ runs over all coupled pairs, and $h_{i}$, $\zeta_{i,j}$ are coefficients in units of angular frequency. $H_{i}$ corresponds to a discrepancy between the actual qubit frequency and our estimated value of it (determined by the precision of the calibration experiment and ensuing drifts), and $V^{zz}_{ij}$ is a residual crosstalk between coupled qubits, resulting from uncontrolled interactions of ZZ type (also referred to as Ising coupling).
In physical systems the crosstalk would often couple nearest-neighbor qubits that would typically be data and parity (auxiliary) qubits, and parity qubits are lacking in our model to begin with. Nevertheless, effective longer-range interaction is also possible, and we consider this coupling as a representative term of coherent dynamics between qubits that are coupled either directly or indirectly. Some 2Q coherent dynamics is also expected when taking into account the over/under rotations of realistic 2Q gates. We note that another prevalent interaction term is the XY (“exchange” or “flip-flip”) coupling [17], which is nondiagonal in the Z basis and introduces some further complexity into the dynamics. For completeness we briefly analyze XY coupling in App. H, strengthening the main conclusions of this work.
@@ III.1 Estimating error correction success rates
The next question is how to extract a useful metric from $\rho(t)$. As discussed further below, an explicit analysis that accounts for all possible syndromes and their corresponding recoveries could become challenging with large codes due to the exponential scaling of their number. Instead, we present in this subsection a simple and efficient approach to compute the code performance, suitable within the code capacity model and somewhat advantageous with larger codes.
Every error correction code has a set of correctable errors that the code can rectify, $C\equiv\left\{C_{j}\right\}$. These errors can be chosen in such a way that their application to a logical code state results in orthogonal states, as discussed in [52]. Starting with an initial ideal logical state $|\psi_{0}\rangle$, we define $|\varphi_{j}\rangle=C_{j}|\psi_{0}\rangle$. Following the noise operation (before the syndrome extraction and recovery process) the components that overlap with the states $|\varphi_{j}\rangle$ will be successfully corrected by the error correction procedure. The components of the state that do not overlap with $|\varphi_{j}\rangle$ are outside of the space of correctable errors, and successful error correction is not guaranteed. Therefore, for a given state $\rho$, the probability of guaranteed success is given by
@@
It is worth highlighting the distinction between this method and the conventional approach for assessing the success rate of a code (based on a stochastic Pauli model) presented in Sec. I, in which every time a logical operation is applied is categorized as a failure. The approach that we take here depends on the initial state; if the noise results in a logical operation that does not alter the initial state, for example, a logical $Z_{L}$ applied to the state $|0\rangle_{L}$, it is considered a success. In this sense, the conventional approach is a worst-case approach deeming every logical operator as a logical error, and this may lead to slight discrepancies between the two methods when applied on the same Pauli noise.
@@ III.2 Estimating a pseudo-threshold
A quantum error correction code is advantageous compared to a physical qubit only for noise parameters such that the probability of an error on the physical qubit is larger than the probability of a logical error on the logical qubit with error correction. Assuming a monotonous dependence on one common noise parameter (where the logical qubit is advantageous for small enough values of the noise parameter), the value of this parameter for which the physical and logical error probabilities are equal determines the pseudo-threshold [36, 61]. In our case, this common parameter is the time $t$ between QEC cycles.
We compute the infidelity on a physical qubit by assuming a single physical qubit standing idle and suffering from the 1Q noise terms for time $t$. Since the 2Q crosstalk does not affect a single isolated physical qubit when standing idle, defining the pseudo-threshold on a single qubit possibly gives the physical qubit an advantage. We note that the standard way to compute the pseudo-threshold must account for the noise affecting logical states during the preparation, stabilizer measurements and corrections. However, here we are within the simplified framework where all of these are noiseless, giving the logical qubit an advantage, so the pseudo-threshold we obtain is merely a code-capacity pseudo-threshold.
@@ III.3 The logical density matrix
To summarize, $\alpha$ is the infidelity of the corrected state with the ideal logical state, or in other words the probability of failure of the correction, and $\beta$ is a measure of coherent logical error that accumulates within the code space. We note that $\eta$ which was studied extensively above, is an approximation of $\alpha$ without explicitly performing the corrections in the circuit. The difference between them stems from eta including only the projection on correctable errors, while alpha is computed on the corrected state, and so could, in general, contain contributions also from non-correctable errors, if following the syndrome projection and correction the resulting state has an overlap with the initial state. Within the code-capacity model, when the projection onto the syndrome subspaces and the correction are both ideal, $\rho^{P}$ is guaranteed to be within the code space.
@@ IV.3 The Pauli approximation
The main advantage of a Pauli channel is that it allows one to compute the syndromes without simulating the state evolution but rather with the parity matrix, as explained in Sec. I. The main physically-driven motivations behind this approximation stems from the influence of the syndrome measurement on the noise channel, which reduces many non-Pauli terms, explored in more detail in App. F.
The most precise way to perform the Pauli approximation is by approximating the whole-system quantum channel. However, this requires constructing the channel which may be numerically as challenging as simulating its action on particular states (or even more). Here we choose a less exact but easier to compute approximation, which consists of approximating each individual channel separately. The channels $\mathcal{K}_{\text{1Q}}$ and $\mathcal{U}^{zz}_{\text{2Q}}$ are each approximated by 1Q and 2Q Pauli channels, respectively, which are then composed. A more detailed explanation on the different possible ways to perform this Pauli approximation is given in App. G. We note that the alternative method of approximating the 1Q and 2Q channels together into one 2Q Pauli channel is not a better choice due to the complication arising from the need to “split” the 1Q channels into the different 2Q channels in which each qubit is included, which reduces the accuracy of the approximation.
@@ V.2 The long time limit
$t$ [$\mu{\rm s}$] Dynamical Composite, order 1 Composite, order 2 Pauli
@@
[…]
For the Kraus channels in the long-time limit, we have $\rho(t\to\infty)\to|0\rangle\langle 0|^{\otimes N}$ for any finite value of $T_{1}<\infty$. However, for the Pauli channels in the long-time limit, we have $\rho(t\to\infty)\to 1\kern-2.5pt\text{l}/2^{N}$, since the 1Q Pauli approximation reduces to the fully depolarizing channel in this limit. This difference in the steady state necessarily implies that at least for $t\gtrsim T_{1}$ (if not earlier), we can a-priori expect any Pauli approximation to break  [^3] The steady state being the fully mixed state results in $\eta\to 1/2$ in the Pauli approximations for any initial state. In contrast, noticing that the state $|0\rangle_{L}$ ($|1\rangle_{L}$) is a superposition over the 5-bit strings with even (odd) number of 1’s, we can calculate the fidelity and correctability of the state in the composite-channel approximation at $t\to\infty$, which gives us $\eta\to 3/8$ for $|0\rangle_{L}$, $\eta\to 5/8$ for $|1\rangle_{L}$ and $\eta\to 1/2$ for $|+\rangle_{L}$..
@@ V.3 Dominant ZZ crosstalk noise
This data strengthens the observations made above, that for the studied parameters the composite-channel approximation is a very good approximation up to intermediate times, while the Pauli approximation works well only for relatively short time-scales. Regarding the composite-channel approximation losing accuracy, this results from the noncommutativity of the 2Q crosstalk channel with amplitude damping. The relevant parameter is $|\zeta|t\times t/T_{1}$ that remains a relatively small parameter (reaching $\approx 0.28$ for $t=15\,\mu$s at the right boundary of Fig. 2). This is a general observation about the nature of the noise model, and the effect of noncommuting terms in the approximation is further explored in App. I.
In the current example, the 2Q noise is the largest term, and is one that cannot be corrected with the employed standard (1Q-noise) decoder of the 5Q code. This results in a near-parabolic dependence on $t$ in Fig. 2(a) consistent with Eqs. (34)-(35). The Pauli approximation overestimates the failure rate when leaving the short time regime, plausibly since it does not describe the coherent dynamics accurately enough (when in this timescale, coherently accumulated phases can lead to cancellations of errors that are impossible in the Pauli description). In order to gain more insight into this, we consider in the next subsection a decoder tailored to the noise.
@@ V.4 Modified decoders
In Fig. 3 we present the difference between $\eta$ with the FC ZZ decoder and with the standard decoder, showing that the FC ZZ decoder is beneficial in a wide range of experimentally relevant parameters. Specifically, the parameter values of Eqs. (40)-(41) fall in the regime where the modified decoder is advantageous. Therefore, applying the FC ZZ decoder to the dynamics with the same parameters as in Fig. 2(a), the FC ZZ decoder performs much better than the regular decoder, obtaining much lower values of $\eta$, as can be seen in Fig. 4. Since Z and ZZ errors are corrected at the leading order, the dominant errors are single-qubit flips ($T_{1}$), with a linear time dependence as in Eq. (30). The Pauli approximation begins to diverge from the dynamical simulation at $t\approx 0.7\mu$s, underestimating the failure rate. In other words, even at lower failure rates and when the code is adapted to the dominant noise terms, the Pauli approximation could lose its accuracy already in the short time regime, and not provide even an upper bound on the logical error.